OpenStudy (delirium):

A savings account with a current balance of $5,000 earns 2.75% annual interest, compounded continuously. To the nearest dollar, what will be the account balance in 10 years?

OpenStudy (delirium):

My answer: A=P(1+rt) A=5,000(1+2.75*10) A=5,000(1+27.5) A=5,000(28.5) A=142,500

OpenStudy (delirium):

But I'm hyper-sensitive to getting anything wrong at this point, as my grade has dropped to a 90.78%, and I need to keep it in the 90s. Only six more assignments to go, but I'm in danger of blowing it. :S

OpenStudy (delirium):

They also tell me to use the following formula, which makes NO sense...I mean, what does 'n' represent??